OpenStudy (anonymous):

what is the function inverse of f(x)= 3 (squareroot of x)

OpenStudy (anonymous):

\[y=3\sqrt{x}\] switch \(x\) and \(y\) \[x=3\sqrt{y}\] solve for \(y\) \[\frac{x}{3}=\sqrt{y}\] \[\left( \frac{x}{3}\right)^2=y\] \[f^{-1}(x)=\frac{x^2}{9}\]

OpenStudy (anonymous):

unless of course you meant \[y=\sqrt[3]{x}\] which is a whole different story

OpenStudy (anonymous):

thank you ! and nope thats correct

OpenStudy (anonymous):

yw